# Tumor-anchored deep feature random forests for out-of-distribution detection in lung cancer segmentation

### v1

@@ 3.2 Scan-level OOD detection approach
Ou OOD detection pipeline consists of the following four steps (Figure 2):
  • 1.
    Regions of interest extraction: The task-specific segmentation model (Subsection 3.1) was used to generate segmentation of tumors from the entire scan. Detected tumor regions from each scan were used to extract multiple 3D ROIs, ensuring the tumor appears within each ROI volume without precise centering (Figure 2a).
  • 2.
    Feature extraction. We repurposed the fine-tuned backbone encoder from the segmentation model to extract multi-scale features from the five encoder stages: the patch embedding layer (48 channels) and the four Swin Transformer stages (96, 192, 384, and 768 channels, respectively). All features were aggregated into a single scan-level descriptor through global average pooling (GAP) describing the scan as a vector of features (Figure 2b).
  • 3.
    Training the OOD detector. Feature vectors were extracted from ID lung cancer datasets as well as a representative set of OOD scans from tumor anchored regions as described above and used to train the RF classifier for each cohort once. This approach enabled our detector to learn subtle decision boundaries that distinguish shifts even between different diseases affecting the same anatomic region (Figure 2c).
  • 4.
@@ 4.1 Implementation details
All models were implemented in PyTorch (paszke2019pytorch) and MONAI (cardoso2022monai). All scans were intensity-normalized using the lung window (Hounsfield units (HU): [400,400][-400,400]) and resampled to a uniform voxel spacing of $1~\text{mm}^{3}$.
Segmentation. Models were fine-tuned using cross-entropy and Dice loss with a batch size of 16 distributed across four NVIDIA GPUs, with an input size of 128 $\times$ 128 $\times$ 128 voxels. We used a learning rate of $2\times 10^{-4}$ with linear warm-up (50 epochs) and cosine annealing over 1,000 epochs. Data augmentations included random flips, rotations, affine transformations, and illumination adjustments, consistent with tang2022self; jiang2022self. For inference on the full scan volume, we used a sliding window with 50% Gaussian overlap to obtain the segmentations.
OOD Detection. Random forest classifiers employed 1,000 trees with a maximum depth of 20 and balanced class weights (using default scikit-learn hyperparameters). We extracted deep feature representations from $n=4$ tumor-anchored 3D ROIs per scan for all ID and OOD datasets. All ROIs were used individually during training to increase training samples, and probabilities from all ROIs were averaged at the scan level as an ensemble for inference. For MaxSoftmax, MaxLogit, and energy, scan-level scores were obtained by mean-aggregating tumor (positive class) predictions across all voxels using the corresponding metric.
@@ 4.2 Datasets
Backbone pretraining. The backbone encoder was pretrained using the self-distilled masked image transformer (SMIT) framework on 10,432 3D CT scans, encompassing public and institutional datasets, following jiang2022self; tang2022self; jiang2024self.
Comparison to other pretraining strategies and architectures. Additionally, we pretrained using SimMIM (xie2022simmim), iBOT (zhou2021ibot), and Swin UNETR (tang2022self) to assess the impact of pretraining strategy on both accuracy and OOD robustness. To analyze the effect of model depth, we pretrained a lite backbone configuration (22222-2-2-2) using SMIT. Finally, we evaluated the publicly available Swin UNETR pretrained checkpoint (marked as ) to assess the impact of pretraining data.
Fine-tuning. We combined the pretrained backbone as the encoder with a randomly initialized U-Net decoder and fine-tuned using $N=317$ 3D scans from publicly available non-small cell lung cancer (NSCLC) Radiomics dataset (aerts2015data) with radiologist-provided tumor delineations.
Testing. We used the NSCLC Radiogenomics dataset (bakr2017data), consisting of 140 images from various scanners (GE, Siemens), as the ID test set. OOD evaluation used 1,916 3D CT scans from four public datasets: pulmonary embolism (RSNA PE, N=1,225N=1,225(colak2021rsna), COVID-19 negative patients (MIDRC C19, N=120N=120(ricord2021covid), kidney cancer (KiTS, N=489N=489(heller2023kits21), and healthy scans of the pancreas (PancreasCT, N=82N=82(roth2015deeporgan). The near-OOD datasets (RSNA PE, MIDRC C19) consist of chest CTs involving non-cancerous pathologies such as pulmonary embolism and inflammation, depicting appearances similar to lung lesions. The far-OOD datasets (KiTS, PancreasCT) include cases with larger field-of-view than training with healthy organs and non-lung cancers.
@@ 4.3 Evaluation metrics
Segmentation accuracy on ID datasets was measured using the Dice similarity coefficient (DSC) and the 95th percentile Hausdorff distance (HD95). OOD detection accuracy was computed using the area under the receiver operating characteristic curve (AUROC) and false positive rate at 95% true positive rate (FPR at 95%TPR, or FPR95). AUROC quantifies the separability between ID and OOD cases (davis2006relationship), with higher values indicating better discrimination and FPR95 measures the proportion of OOD cases misclassified as ID when sensitivity is fixed at 95% (liang2017enhancing), where lower values are better.
To ensure meaningful AUROC computation, the number of OOD scans was balanced with ID test cases ($N=140$) by sampling OOD examples with replacement in datasets (RSNA PE and KiTS) with a much larger number of OOD scans and averaging the computed metrics across multiple draws. Splits ensured zero patient overlap with consistent OOD case counts across runs. This evaluation approach avoids biases from imbalanced evaluation sets while ensuring stable ID-OOD ratios, following best practices for robust OOD evaluation (szyc2023out; humblot2023beyond).
For segmentation performance, we reported the mean and standard deviation across test cases. For OOD detection performance, we reported the point estimates with 95% bootstrap confidence intervals computed over 100 runs with matched random seeds to reflect sampling variability across different train-validation splits.
@@ 4.4 Comparative OOD detection methods
Baseline comparisons included commonly used OOD detection methods, namely, MaxSoftmax (hendrycks17baseline), MaxLogit (hendrycks2022logits), and energy (liu2020energy), performed by computing scan-level scores by mean-aggregating the tumor (positive class) predictions across all voxels (see Appendix B).
In addition, a RF classifier was created using handcrafted radiomics features ($N=293$) called RF-Radiomics, computed within the segmented tumor regions following the Image Biomarker Standardization Initiative (IBSI) guidelines (zwanenburg2020image) with open-source software PyCERR (apte2018extension). Details of individual radiomic features are in Table 5. The high correlation among radiomic features was addressed by employing recursive feature elimination to select a more compact and informative set of radiomic features. Of note, recursive feature elimination was not used for training the RF-Deep classifier.
Finally, the Mahalanobis distance-based OOD detector (MD-Deep) (lee2018simple) was computed with respect to deep features of the ID data, modeled as a single Gaussian in feature space using Ledoit-Wolf covariance estimation. Hence, MD-Deep computes OOD scores as the distance of the each scan with respect to the fitted ID distribution. We excluded ODIN (liang2017enhancing), as it requires temperature scaling and gradient perturbations on fine-tuning data, which in real clinical settings is often unavailable when applied to new OOD cohorts.
@@ 5 Results
5.1 Segmentation performance on ID datasets
Segmentation performance for the various analyzed pretraining strategies is shown in Figure 3(a) for the ID testing dataset. All pretrained methods were similarly accurate, with the lowest accuracies produced by iBOT, followed by the publicly available Swin UNETR checkpoint.
@@ 5.1 Segmentation performance on ID datasets
Overall, the SMIT-pretrained model was the most accurate, followed by the Swin UNETR method, and was thus selected for benchmarking the OOD detection method using RF-Deep througout this paper.
@@
5.2 Out-of-distribution detection performance
RF-Deep is the most accurate method for both near-OOD and far-OOD datasets (Table 1). On the near-OOD datasets (RSNA PE and MIDRC C19), RF-Deep achieved AUROCs of 95.80 and 93.30, outperforming the next best method (MaxLogit) by approximately 7.7% and 4.0%, respectively. Similarly, RF-Deep reduced FPR95 by 18–25 absolute points compared to the next closest competitor. These improvements remain consistent within the 95% confidence intervals, with non-overlapping intervals confirming statistically significant superiority over all baseline methods, highlighting the stability of RF-Deep across bootstrap resampling.
RF-Deep also achieved near-perfect separability on the far-OOD datasets using KiTS (AUROC 100, FPR95 0.10%) for kidney cancers and PancreasCT (AUROC 100, FPR95 0.00%) containing healthy abdominal scans, effectively eliminating false alarms for anatomically distant pathologies. Although it was less accurate on the near-OOD cohorts compared to its accuracy on the far-OOD cohorts, the gap in AUROC (RSNA PE: ++3.5%, MIDRC C19: ++6.7%) was the smallest compared to all other methods, with the next smallest gap observed for RF-Radiomics (RSNA PE: ++7.5%, MIDRC C19: ++4.9%).
RF-Radiomics, while competitive on the far-OOD datasets (AUROC of 96.20 on KiTS and 96.10 on PancreasCT with relatively low FPR95), produced moderate error rates on the near-OOD cohorts (AUROCs 88.70 and 91.30, FPR95s around 3041%), indicating that handcrafted features are less effective than deep features at capturing subtle differences in the appearance of pathologies (analyzed further in Sections 5.3.1 and 5.3.2).
MD-Deep, although highly accurate for detection on far-OOD datasets (AUROC 99.2 on both KiTS and PancreasCT) was less effective and similar to logit-based methods on near-OOD cohorts (AUROC 87.3 on RSNA PE and 84.5 on MIDRC C19, with FPR95 around 45-59%). MD-Deep was also less accurate than RF-Deep on near-OOD datasets, despite both methods using identical encoder features, indicating that feature selection performed by the RF was essential to achieve better separation of ID from OOD scans. Additional evaluation to assess if feature reduction enhanced performance showed no improvement compared to the standard MD-Deep approach (Figure 11).
@@ 5.2.1 Impact of model pretraining on OOD detection
Next, we analyzed whether the choice of model pretraining, model size (SMIT versus SMIT Lite), and pretraining data (Swin UNETR versus Swin UNETR) impacted the accuracy of RF-Deep. RF-Deep remained the most accurate method across all backbone architectures (Table 6), consistently achieving near-perfect accuracy on the far-OOD datasets. Table 7 further demonstrated that RF-Deep’s superiority over logit-based and radiomics methods was consistent across all tested backbones, with RF-Deep outperforming the next best method by large margins on the near-OOD datasets. Accuracy differences between the SMIT and SMIT Lite models on the near-OOD cohorts were modest (RSNA PE AUROC: 93.7 vs 94.1; MIDRC C19 AUROC: 93.2 vs 90.4; RSNA PE FPR95: 17.6% vs 20.4%; MIDRC C19 FPR95: 26.5% vs 30.5%), with SMIT Lite showing slightly better performance on RSNA PE but marginally lower accuracy on MIDRC C19.
@@ 5.3.1 RF-Deep extracts features that are better separated than RF-Radiomics
Figure 4 visualizes the t-SNE embeddings (perplexity=30, 1,000 iterations) of deep features (panel a) and radiomics features (panel b) for the ID and OOD datasets collectively. RF-Deep features exhibit better cluster separation between the ID and OOD datasets, with minimal overlap between ID and anatomically distinct far-OOD datasets (KiTS, PancreasCT) and moderate overlap with near-OOD datasets (RSNA PE, MIDRC C19), consistent with their increased detection difficulty. Radiomics features produced poorly separated clusters for the various datasets particularly for the near-OOD datasets with respect to the ID cases, explaining the relatively lower accuracies (Table 1), and less compact OOD score distributions than RF-Deep between ID and OOD datasets as shown in Figure 5.
(a) Radiomics features
(b) Deep features
Figure 5: Scan-level OOD detection score distributions for RF-Radiomics (a) and RF-Deep (b) across 100 matched-seed runs.

### v2

@@ 3.2 Scan-level OOD detection approach
Our OOD detection pipeline consists of the following four steps (Figure 2):
  • Regions of interest extraction: The task-specific segmentation model (Subsection 3.1) was used to generate tumor segmentations from the entire scan. Detected tumor regions were used to extract multiple 3D ROIs via random spatial offsets, ensuring the tumor appears within each ROI without precise centering (Figure 2a).
  • Feature extraction. We repurposed the fine-tuned backbone encoder from the segmentation model to extract multi-scale features from the five encoder stages: the patch embedding layer (48 channels) and the four Swin Transformer stages (96, 192, 384, and 768 channels, respectively). All features were aggregated into a single scan-level descriptor through global average pooling (GAP), producing a single feature vector per scan (Figure 2b).
  • Training the OOD detector. Feature vectors were extracted from ID lung cancer datasets as well as a representative set of OOD scans from tumor-anchored regions as described above and used to train the RF classifier. This approach enabled our detector to learn subtle decision boundaries that distinguish shifts even between different diseases affecting the same anatomic region (Figure 2c).
@@ 4.1 Implementation details
All models were implemented in PyTorch (Paszke et al., 2019) and MONAI (Cardoso et al., 2022). Our method operates on full 3D CT volumes at full resolution, resampled to a uniform isotropic voxel spacing of $1~\text{mm}^{3}$. All scans were clipped to the lung window (Hounsfield units (HU): [400,400][-400,400]) and then intensity-normalized. During inference, the segmentation model processes the entire volume using sliding window inference ($128\times 128\times 128$ voxel patches with 50% Gaussian overlap) to generate tumor segmentations. For OOD detection, every detected tumor seeds multiple tumor-anchored 3D ROIs ($128\times 128\times 128$ voxels each) around the predicted tumor areas; the OOD scores from these regions are combined to generate a scan-level prediction.
Segmentation. Models were fine-tuned using cross-entropy and Dice loss with a batch size of 16 distributed across four NVIDIA GPUs, with an input size of 128 $\times$ 128 $\times$ 128 voxels. We used AdamW optimizer (Loshchilov and Hutter, 2017) with a learning rate of $2\times 10^{-4}$, linear warm-up (50 epochs), and cosine annealing over 1,000 epochs. Data augmentations included random flips, rotations, affine transformations, and illumination adjustments, consistent with Tang et al. (2022); Jiang et al. (2022).
OOD Detection. Random forest classifiers employed 1,000 trees with a maximum depth of 20 (using default scikit-learn hyperparameters). We extracted deep feature representations from $n=4$ tumor-anchored 3D ROIs per scan for all ID and OOD datasets. All ROIs were used individually during training to increase training samples, and probabilities from all ROIs were averaged at the scan level as an ensemble for inference. For MaxSoftmax, MaxLogit, and energy, scan-level scores were obtained by mean-aggregating tumor (positive class) predictions across all voxels using the corresponding metric.
@@ 4.2 Datasets
Backbone pretraining. The backbone encoder was pretrained using the self-distilled masked image transformer (SMIT) framework on 10,432 3D CT scans, encompassing public and institutional datasets, following Jiang et al. (2022); Tang et al. (2022); Jiang et al. (2024).
Comparison to other pretraining strategies and architectures. Additionally, we pretrained using SimMIM (Xie et al., 2022), iBOT (Zhou et al., 2022), and Swin UNETR (Tang et al., 2022) to assess the impact of pretraining strategy on both accuracy and OOD robustness. To analyze the effect of model depth, we pretrained a lite backbone configuration (22222-2-2-2) using SMIT. Finally, we evaluated the publicly available Swin UNETR pretrained checkpoint (marked as ) to assess the impact of pretraining data.
We do not consider interactive foundation models for segmentation, such as SAM (Kirillov et al., 2023) or its medical variants (e.g., MedSAM (Ma et al., 2024) and VISTA3D He et al. (2025)), as these methods are primarily designed for prompt-guided segmentation and require user-provided inputs (points or bounding boxes), which precludes fully automated deployment. Additionally, we exclude self-supervised 2D models such as DINOv2 (Oquab et al., 2023), which are pretrained on natural images and operate on individual slices, thereby losing critical inter-slice spatial context necessary for distinguishing 3D pathological structures in CT volumes (that is, relying on 2.5D processing).
Fine-tuning. We combined the pretrained backbone as the encoder with a randomly initialized U-Net decoder and fine-tuned using $N=317$ 3D scans from the publicly available non-small cell lung cancer (NSCLC) Radiomics dataset (Aerts et al., 2015) with radiologist-provided tumor delineations. The original dataset contains 422 scans; 317 were selected based on radiologist assessment and verification of tumor boundary quality.
Testing. We used the NSCLC Radiogenomics dataset (Bakr et al., 2017) as the ID test set. From the original 211 scans, $N=140$ scans from various scanners (GE, Siemens) were selected based on radiologist assessment and verification of tumor boundaries and image quality. The ID test set is completely separate from the training set, acquired from different patients and lung cancer stages, providing external validation. OOD evaluation used 1,916 3D CT scans from four public datasets: pulmonary embolism (RSNA PE, $N=1,225$ selected from 3,114 scans, preserving the marginal distribution of labels in the full PE-positive dataset) (Colak et al., 2021), COVID-19 negative patients (MIDRC C19, $N=120$, all used) (Tsai et al., 2021), kidney cancer (KiTS, $N=489$, all used) (Heller et al., 2023), and healthy scans of the pancreas (PancreasCT, $N=82$, all used) (Roth et al., 2015). The near-OOD datasets (RSNA PE, MIDRC C19) consist of chest CTs involving non-cancerous pathologies such as pulmonary embolism and inflammation, depicting appearances similar to lung lesions. The far-OOD datasets (KiTS, PancreasCT) include scans with larger field-of-view than training with healthy organs and non-lung cancers. Patient-level splits ensured zero overlap between training and evaluation; OOD counts were held constant across runs.
@@ 4.3 Evaluation metrics
Segmentation accuracy on ID datasets was measured using the Dice similarity coefficient (DSC) and the 95th percentile Hausdorff distance (HD95). OOD detection accuracy was computed using the area under the receiver operating characteristic curve (AUROC) and false positive rate at 95% true positive rate (FPR at 95%TPR, or FPR95). AUROC quantifies the separability between ID and OOD scans (Davis and Goadrich, 2006), with higher values indicating better discrimination. FPR95 measures the proportion of OOD scans misclassified as ID when sensitivity is fixed at 95% (Liang et al., 2018), where lower values are better.
To ensure meaningful AUROC computation, the number of OOD scans was balanced with ID test scans ($N=140$) by subsampling OOD examples without replacement from larger datasets (RSNA PE and KiTS) and averaging the computed metrics. Splits ensured zero patient overlap with consistent OOD scan counts. This evaluation approach avoids biases from imbalanced evaluation sets while ensuring stable ID-OOD ratios, following best practices for robust OOD evaluation (Szyc et al., 2023; Humblot-Renaux et al., 2023).
For segmentation performance, we reported the mean and standard deviation across test scans. For OOD detection performance, we reported the point estimates with 95% bootstrap confidence intervals computed over 100 runs with matched random seeds to reflect sampling variability across different train-validation splits.
@@ 4.4 Comparative OOD detection methods
Baseline comparisons included commonly used OOD detection methods, namely, MaxSoftmax (Hendrycks and Gimpel, 2017), MaxLogit (Hendrycks et al., 2022), and energy (Liu et al., 2020), performed by computing scan-level scores by mean-aggregating the tumor (positive class) predictions across all voxels (see Appendix B).
In addition, a RF classifier was created using handcrafted radiomics features ($N=293$) called RF-Radiomics, computed within the segmented tumor regions following the Image Biomarker Standardization Initiative (IBSI) guidelines (Zwanenburg et al., 2020) with open-source software PyCERR (Apte et al., 2018). Critically, RF-Radiomics uses identical outlier exposure as RF-Deep (same ID and OOD training splits), isolating the effect of feature representation: any performance difference between RF-Deep and RF-Radiomics is attributable to deep features versus handcrafted features, not to differential access to OOD supervision. Details of individual radiomic features are in Table 5. The high correlation among radiomic features was addressed by employing recursive feature elimination to select a more compact and informative set of radiomic features. Of note, recursive feature elimination was not used for training the RF-Deep classifier.
Finally, the Mahalanobis distance-based OOD detector (MD-Deep) (Lee et al., 2018) was computed with respect to deep features of the ID data, modeled as a single Gaussian in feature space using Ledoit-Wolf covariance estimation. Hence, MD-Deep computes OOD scores as the distance of each scan with respect to the fitted ID distribution. MD-Deep uses identical backbone features as RF-Deep but operates without outlier exposure, isolating the contribution of OE and the RF classifier’s learned decision boundaries. Comparing RF-Deep to MD-Deep thus reveals the benefit of outlier exposure and non-linear classification over purely unsupervised density estimation on the same feature space. We excluded ODIN (Liang et al., 2018) because it relies on temperature scaling and input perturbations calibrated on fine-tuning data, parameters that are often unavailable when deploying models to new out-of-distribution clinical datasets.
@@ 5 Results
5.1 Segmentation performance on ID dataset
Segmentation performance for the various analyzed pretraining strategies is shown in Figure 3(a) for the ID testing dataset. All pretrained methods were similarly accurate, with iBOT producing the lowest accuracy, followed by the publicly available Swin UNETR checkpoint.
@@ 5.1 Segmentation performance on ID dataset
Overall, the SMIT-pretrained model was the most accurate, followed by the Swin UNETR method, and was thus selected for benchmarking the OOD detection method using RF-Deep throughout this paper.
5.2 Out-of-distribution detection performance
RF-Deep is the most accurate method for both near-OOD and far-OOD datasets (Table 1). On the near-OOD datasets (RSNA PE and MIDRC C19), RF-Deep achieved AUROCs of 95.80 and 93.30, outperforming the next best method (MaxLogit) by approximately 7.0 and 4.0 percentage points, respectively. Similarly, RF-Deep reduced FPR95 by 18–25 absolute points compared to the next closest competitor. These improvements remain consistent within the 95% confidence intervals, with non-overlapping intervals confirming statistically significant superiority over all baseline methods, highlighting the stability of RF-Deep across bootstrap resampling.
@@ 5.2 Out-of-distribution detection performance
RF-Deep also achieved near-perfect separability on the far-OOD datasets using KiTS (AUROC 100, FPR95 0.10%) for kidney cancers and PancreasCT (AUROC 100, FPR95 0.00%) containing healthy abdominal scans, effectively eliminating false alarms for anatomically distant pathologies. Although it was less accurate on the near-OOD datasets compared to its accuracy on the far-OOD datasets, the gap in AUROC (RSNA PE: ++3.5%, MIDRC C19: ++6.7%) was the smallest compared to all other methods, with the next smallest gap observed for RF-Radiomics (RSNA PE: ++7.5%, MIDRC C19: ++4.9%).
RF-Radiomics, while competitive on the far-OOD datasets (AUROC of 96.20 on KiTS and 96.10 on PancreasCT with relatively low FPR95), produced moderate error rates on the near-OOD datasets (AUROC 88.70 and 91.30, FPR95 3041%), indicating that handcrafted features are less effective than deep features at capturing subtle differences in the appearance of pathologies (analyzed further in Sections 5.3.1 and 5.3.2).
MD-Deep, although highly accurate for detection on far-OOD datasets (AUROC 99.2 on both KiTS and PancreasCT) was less effective and similar to logit-based methods on near-OOD datasets (AUROC 87.3 on RSNA PE and 84.5 on MIDRC C19, with FPR95 around 45-59%). MD-Deep was also less accurate than RF-Deep on near-OOD datasets, despite both methods using identical encoder features, indicating that feature selection performed by the RF was essential to achieve better separation of ID from OOD scans. Additional evaluation to assess if feature reduction enhanced performance showed no improvement compared to the standard MD-Deep approach (Figure 10).
@@
The controlled comparisons isolate specific contributions: (1) RF-Deep vs. RF-Radiomics (same classifier, same OE, different features) demonstrates that deep features outperform handcrafted radiomics, particularly on near-OOD where subtle pathological differences matter; (2) RF-Deep vs. MD-Deep (same features, different classifier, with vs. without OE) demonstrates that outlier exposure combined with RF’s non-linear decision boundaries outperforms unsupervised density estimation; (3) RF-Deep vs. logit-based methods (different paradigm entirely) demonstrates that encoder features are more discriminative than output-layer uncertainties for OOD detection.
@@ 5.2.1 Impact of model pretraining on OOD detection
Next, we analyzed whether the choice of model pretraining, model size (SMIT versus SMIT Lite), and pretraining data (Swin UNETR versus Swin UNETR) impacted the accuracy of RF-Deep. RF-Deep remained the most accurate method across all backbone architectures (Table 6), consistently achieving near-perfect accuracy on the far-OOD datasets. Table 7 further demonstrated that RF-Deep’s superiority over logit-based and radiomics methods was consistent across all tested backbones, with RF-Deep outperforming the next best method by large margins on the near-OOD datasets. Accuracy differences between SMIT and SMIT Lite models on near-OOD datasets were modest, with SMIT Lite showing slightly better performance on RSNA PE but marginally lower on MIDRC C19.
@@ 5.3.1 RF-Deep extracts features that are better separated than RF-Radiomics
Figure 4 visualizes the t-SNE embeddings (perplexity=30, 1,000 iterations; consistent across random seeds) of deep features (panel a) and radiomics features (panel b) for the ID and OOD datasets collectively. RF-Deep features exhibit better cluster separation between the ID and OOD datasets, with minimal overlap between ID and anatomically distinct far-OOD datasets (KiTS, PancreasCT) and moderate overlap with near-OOD datasets (RSNA PE, MIDRC C19), consistent with their increased detection difficulty. Radiomics features produced poorly separated clusters for the various datasets particularly for the near-OOD datasets with respect to the ID scans, explaining the relatively lower accuracies (Table 1).